OpenStudy (anonymous):

i need the plot points for x + y = 14 and x - y = -4.

OpenStudy (anonymous):

x+y = 14 +x-y = -4 2x = 10 x = 5 5-y = -4 -y = -9 y = 9 (x,y) = (5,9)